OpenStudy (cucacula):

if a and b are the roots of the equation 2x²-6x-8=0 , find the value of : a^2 - b^2 .

OpenStudy (amorfide):

is that suppose to be 2x²-6x-8=0

OpenStudy (cucacula):

yes, typo :)

OpenStudy (cucacula):

A²-B²= -5 ?? @amorfide

OpenStudy (cucacula):

A²-B²= -17 ?

OpenStudy (amorfide):

oops wait i got mixed up ignore that i will do your solution now A and b are roots of this equation 2x²-6x-8=0 A+B=-b/a A+B=6/2 A+B=3 AB=c/a AB=-8/2 AB=-4 A²-B²=(A-B)(A+B) (A-B)²=(A+B)²-4AB=3²-4(-4)=9+16=25 so A-B=square root of 25 A-B= plus or minus 5 A²-B²=(A+B)(A-B)=3(5) or A²-B²=(A+B)(A-B)=3(-5) A-b= 15 or -15

OpenStudy (cucacula):

the final asnwer for a = 15 and b = -15 @amorfide

OpenStudy (amorfide):

A²-B²=15 or -15

OpenStudy (cucacula):

thank you so much @amorfide question solved :)